OpenStudy (anonymous):

find a quadratic equation with the following roots . -3 and 4

OpenStudy (anonymous):

(x + 3)(x - 4) = 0 x^2 - x - 12 = 0

OpenStudy (anonymous):

x+3=0 or x-4=0 x^2 -x - 12 = 0

OpenStudy (anonymous):

when we say a quadratic equation has roots a and b, we mean that the values a and b "satisfy" the equation. in other words, if you substitute x with either a or b, the resulting value of the function is 0. this is why roots are also called zeros

OpenStudy (anonymous):

(x+3)(x-4)=x^2-x-12

OpenStudy (anonymous):

sum and product....x^2-s.or.x+p.or.......sum is 1 and product is -12